Bilt Rite, Inc. has sales of $610,000. The cost of goods sold is equal to 70% of sales. The beginning accounts receivable balance is $21,000 and the ending accounts receivable balance is $25,000. How long on average does it take the firm to collect its receivables?  

26.52 days

21.07 days

14.09 days

25.98 days

13.76 days
Answer #1

Average AR=(Beginning AR+Ending AR)/2

=(21000+25000)/2=$23000

Average days to collect receivables=(Average AR/Sales)*365 days

=($23000/610,000)*365 days

=13.76 days(Approx).
Answer #2

To calculate the average collection period, we need to divide the accounts receivable turnover by the number of days in the year.

The accounts receivable turnover is calculated as:

Accounts Receivable Turnover = Sales / Average Accounts Receivable

To find the average accounts receivable, we use the formula:

Average Accounts Receivable = (Beginning Accounts Receivable + Ending Accounts Receivable) / 2

Using the values given in the problem, we can calculate:

Cost of Goods Sold = 70% * Sales Cost of Goods Sold = 0.7 * $610,000 Cost of Goods Sold = $427,000

Average Accounts Receivable = ($21,000 + $25,000) / 2 Average Accounts Receivable = $23,000

Accounts Receivable Turnover = $610,000 / $23,000 Accounts Receivable Turnover = 26.52

Assuming a 365-day year, the average collection period can be calculated as:

Average Collection Period = 365 days / Accounts Receivable Turnover Average Collection Period = 365 days / 26.52 Average Collection Period = 13.76 days

Therefore, the answer is 13.76 days.

answered by: Hydra Master